# Language Model Alignment in Multilingual Trolley Problems

### v4

@@ Abstract
We evaluate the moral alignment of large language models (LLMs) with human preferences in multilingual trolley problems. Building on the Moral Machine experiment, which captures over 40 million human judgments across 200+ countries, we develop a cross-lingual corpus of moral dilemma vignettes in over 100 languages called MultiTP. This dataset enables the assessment of LLMs’ decision-making processes in diverse linguistic contexts. Our analysis explores the alignment of 19 different LLMs with human judgments, capturing preferences across six moral dimensions: species, gender, fitness, status, age, and the number of lives involved. By correlating these preferences with the demographic distribution of language speakers and examining the consistency of LLM responses to various prompt paraphrasings, our findings provide insights into cross-lingual and ethical biases of LLMs and their intersection. We discover significant variance in alignment across languages, challenging the assumption of uniform moral reasoning in AI systems and highlighting the importance of incorporating diverse perspectives in AI ethics. The results underscore the need for further research on the integration of multilingual dimensions in responsible AI research to ensure fair and equitable AI interactions worldwide. [^1] Our code and data are at https://github.com/causalNLP/moralmachine.
@@ 1 Introduction
The increasingly impressive performance of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Bubeck et al., 2023) also brings safety concerns. First, do LLMs align with human preferences Bengio et al. (2023); Hendrycks et al. (2021b); Anwar et al. (2024)? Second, whose preferences are they most aligned with (Sorensen et al., 2024c)? Prior work on evaluating alignment often takes one of two forms. The first requires continuous real-time evaluation, such as in ChatArena Chiang et al. (2024), where users select which of two responses they prefer. The second is traditional benchmarking with a static dataset where the answers are determined by the judgment of an experimenter or crowdsourced from a limited online panel Sap et al. (2019); Forbes et al. (2020); Atari et al. (2023); Jiang et al. (2021b).
@@
The MultiTP dataset includes 97,520 trolley problem scenarios, on which we evaluate 19 different LLMs (open and closed-weight). Our investigation is structured by five research questions, ranging from measuring the overall alignment across all scenarios to whether (mis)alignment correlates with whether a language is high- or low- resource. Our findings reveal that very few LLMs demonstrate overall alignment with human preferences. Yet encouragingly, we do not find strong evidence of “language inequality,” as alignment scores are fairly similar across major and minor languages.
Main contributions. This work (1) presents Multilingual Trolley Problems, a test set enabling thorough inspection of LLM multi-lingual alignment with human preferences, (2) deploys parametric variation in trolley problem queries, to inspect the causal effect of six different preference dimensions, (3) investigates the alignment of 19 LLMs with human preferences through over 100 languages, and (4) finds that most LLMs do not align well with human preferences on trolley problems, but also in the meantime do not demonstrate bias towards low-resource languages.
@@ 3.1.0.0.1 The Original Human Study
We base our study on the extensive set of human preferences collected in Awad et al. (2018), which serves as the ground truth for alignment. This collection includes approximately 40 million responses from individuals in 233 countries. Following the original study’s methodology, we include data from countries with at least 100 respondents, resulting in a dataset that encompasses 130 countries.
@@ 3.2.0.0.2 Multilingual Variation
We initially developed 920 English vignettes by systematically varying scenarios and character combinations across six moral dimensions, ensuring thorough coverage of all possible character interactions. To expand the scope of our study, we translated this dataset into multiple languages. Using the googletrans Python package, we employed Google Translate to convert the English prompts into 107 supported languages. Although this does not cover every language globally, it encompasses a wide variety, including many low-resource languages with relatively few speakers. Instead of using LLMs, whose translation quality is still uncertain, we use Google Translate, as it is widely recognized as a reliable tool for translating English into other languages, especially for less common languages (Costa-jussà et al., 2022; Jiao et al., 2023; Zhu et al., 2023; Peng et al., 2023). To ensure the accuracy of these translations, we manually reviewed a subset of them in several major languages to confirm that the intended meaning of the prompts was preserved. A full list of the languages used in our trolley problem scenarios is provided in Section B.1.
@@ 3.2.0.0.3 Dataset Statistics
We present comprehensive statistics of the MultiTP dataset in Table 2. The dataset comprises 97,520 trolley problem vignettes, with 920 vignettes for each of the 107 languages. We also provide a snapshot of statistics for four representative languages from different global regions: global west, east, and south. On average, vignettes contain 51 words; English scenarios average 47 words, while Chinese scenarios, due to the linguistic structure, average 78 words per scenario. In contrast, Swahili vignettes are shorter, averaging 38 words. The entire dataset incorporates 6,492 unique words, with a type-token ratio of 0.0013.
@@ 4.0.0.0.2 Preference Assessment
Our test is designed to include six types of systematic variations, allowing us to represent a model’s moral preference with six values: $p_{\text{species}},p_{\text{gender}},p_{\text{fitness}},p_{\text{status}},p_{\text{age}},p_{\text{number}}$. For each dimension $p_{i}$, we report the percentage $p_{i}\in[0,1]$ of the time when a default value prevails, namely sparing humans (over pets), sparing more lives (over fewer lives), sparing women (over men), sparing the young (over the elderly), sparing the fit (over the less fit), and sparing those with higher social status (over lower social status). For example, if a model’s $p_{\text{species}}=1$, it consistently prefers humans over pets, 100% of the time. We post-process LLM responses to extract their binary choices for each query.
@@ 4.0.0.0.3 Misalignment Metric
Calculating misalignment presents a significant challenge due to the different bases of recording human and LLM preferences — by country for humans and by language for LLMs. While in many instances a language corresponds to a country, such as Italian to Italy and Romanian to Romania, there are numerous countries where multiple languages are spoken. In these cases, we compute a weighted average of the misalignment scores for all languages spoken within the country, using the number of speakers per language as weights. We source our language population statistics from Wikipedia and detail our method for mapping languages to countries in Section B.2. We also discuss potential limitations of this approach in Section 6.
@@ 5 Research Questions and Results
Our study explores five research questions (RQs) that examine various aspects of the moral alignment of large language models (LLMs) with human preferences. We begin by assessing the global alignment of LLM preferences with human preferences (RQ1). Next, we analyze how LLMs respond to the six principal dimensions studied in the Moral Machine trolley problems (RQ2). We then investigate if LLMs’ responses vary significantly across different languages and identify clusters of language groups where LLMs exhibit similar behavior (RQ3). We also test a “language inequality” hypothesis to determine if LLMs are more likely to be aligned with high-resource language than with low-resource languages (RQ4). Finally, we conduct a robustness study to evaluate the consistency of LLM responses to various paraphrasings of the same trolley problem prompt (RQ5).
@@ 5.2 RQ2: What Are LLMs’ Preferences on Each Moral Dimension in Multilingual Trolley Problems?
(a) Human preferences.
(b) Llama 3.1 70B (aligned).
(c) GPT-4o Mini (misaligned).
@@
To further understand how each moral dimension contributes to the overall misalignment, we find strong correlations between the overall misalignment and several moral dimensions, especially gender, age, and fitness. With a p-value of less than 0.001, we find the correlation with gender is 0.87, 0.69 for age, and 0.68 for fitness. In Figure 2(b), we show the distribution of preference scores across these dimensions for both three most aligned models (in green) and three least models (in red). This analysis confirms that misaligned models often exhibit extreme preferences on each moral dimension, such as a higher propensity to protect females, the young, and individuals of higher social status, diverging markedly from the more balanced tendencies observed in human subjects.
Table 3: Language sensitivity (Sens.) scores for all models. Higher scores mean more varied misalignment across languages.
Model Sens.
@@
Figure 4: Distribution of preferences by each moral dimension across languages, using the most-aligned model Llama 3.1 70B. The dashed line is the overall human preference on each dimension. Cluster A: Georgian, Filipino, Maltese, etc. B: German, Italian, Ukrainian, etc. C: English, Finnish, Chinese, etc. D: Hungarian, Kazakh, Uyghur, etc. See Section D.1 for the entire list of languages in each cluster, as well as clustering results for other models such as GPT-3 and GPT-4.
@@ 5.3 RQ3: Does LLMs Behavior Depend on the Language of Trolley Problems?
The language sensitivity scores of 19 LLMs, as documented in Figure 4, show that most models exhibit clear sensitivity across their language-specific responses, with standard deviation values ranging from 14.7 to 24.7 across languages. Exploring further, we identify distinctive patterns within the language clusters using K-Means clustering. In Llama 3.1 70B, we identify four distinct language clusters. For example, Cluster A (e.g., Georgian and Filipino) values animal lives more than the other three clusters, which unanimously always favor humans over animals. Cluster D (e.g., Hungarian and Kazakh) has close-to-zero bias against unfit people, with a 50% possibility favoring the fit over the unfit.
@@ 5.4 RQ4: Are LLMs More Misaligned in Low-Resource Languages?
A considerable amount of recent research indicates that LLMs tend to align more closely with high-resource languages, as noted in various studies (Tanwar et al., 2023; Cahyawijaya et al., 2023; Nguyen et al., 2024; Ryan et al., 2024; Zhao et al., 2024). Despite this trend, we seek to determine if the same alignment preference applies to multilingual trolley problems. An important property of MultiTP is that it uses data that is not found online and hence was not seen during the LLMs’ training. This is because we generate our test data on-demand based on parametric variations and translations that we perform specifically for this study. In this context, our research question investigates whether there is a significant positive correlation between the degree of LLM alignment on trolley problems and how widely-used a language is. We measure the number of speakers per language by extracting data from Wikipedia’s language demographics statistics Wikipedia (2024). Then, we calculate the Pearson correlation coefficient between model misalignment scores and the number of speakers for each language.
@@
Most models do not show a significant correlation between the per-language misalignment score and the number of speakers per language, with correlations close to zero. See the entire table of the correlation coefficients of all models in Section D.2. This is an interesting observation unique to the MultiTP trolley problems, which challenges the necessity of the “language inequality” hypothesis. For example, in Llama 3.1 70B, the misalignment scores for the top five most spoken languages are: Chinese 0.38, Hindi 0.51, English 0.58, Spanish 0.68, and Arabic 0.54. This distribution of correlations is about the same as in less spoken languages among our dataset: Bosnian 0.54, Luxembourgish 0.37, Icelandic 0.58, Maltese 0.57, Malayalam 0.54, and Catalan 0.59. We also visualize a misalignment world map in Section D.3 showcasing that there is no strong correspondence between the development level of a country and its alignment score.
@@ 6 Future Work Directions
Beyond the research questions we have asked and answered above, we highlight several areas for improvement and productive future study.
Exploring Jailbreaking for LLMs: Exploring whether jailbroken models maintain similar moral preferences is important, particularly to measure whether or not they exhibit increased bias towards gender or social status. Additionally, jailbreaking can help circumvent LLM refusals to respond. Due to safety tuning in recent LLMs, a significant challenge in our study is the high refusal rate; LLMs often decline to make decisions involving human lives or offer generic responses considering multiple moral criteria. We found that the refusal rate for earlier models like GPT-3 is relatively low at 12.1%, but it increases substantially in more recent models. When responding to different moral dimensions, the models provide more definitive answers about species and the number of lives saved but frequently avoid addressing sensitive issues such as gender. In our research, we employed a basic technique known as token-forcing to mitigate refusal rates. Future studies could investigate more advanced jailbreaking techniques to further lower these rates and uncover deeper underlying preferences in LLMs.
Extension to Different Modalities and More Variations: In our study, we focus exclusively on a text-only setting to investigate the latest text-based LLMs available at the time of our research. Although the original Moral Machine study included a visual demonstration of each scenario [^2] https://moralmachine.net, we recommend that future research also test the alignment of multi-modal foundation models. Additionally, there is an opportunity for future studies to expand the diversity of characters, scenario setups, and dimensions of variation. This expansion would increase the combinatorial variation of moral dilemmas that can be posed as trolley problems. Another point to consider is that the trolley problem is sometimes criticized for being too narrow or unrealistic Steen (2024). Nonetheless, it remains a widely used tool for studying moral decision-making. For this reason, the current study adopts this foundational scenario while encouraging future research to incorporate additional complexities and connect it to real-world situations.
More Language Support: Human languages are vast and diverse. While our study utilizes the 107 languages supported by high-quality translation services from Google Translate, there is a need for further research into low-resource languages. We recommend that future studies expand their scope to include a broader range of these low-resource languages to gain a more comprehensive understanding LLMs’ moral preferences in diverse linguistic contexts.

### v6

@@ Abstract
We evaluate the moral alignment of LLMs with human preferences in multilingual trolley problems. Building on the Moral Machine experiment, which captures over 40 million human judgments across 200+ countries, we develop a cross-lingual corpus of moral dilemma vignettes in over 100 languages called MultiTP. This dataset enables the assessment of LLMs’ decision-making processes in diverse linguistic contexts. Our analysis explores the alignment of 19 different LLMs with human judgments, capturing preferences across six moral dimensions: species, gender, fitness, status, age, and the number of lives involved. By correlating these preferences with the demographic distribution of language speakers and examining the consistency of LLM responses to various prompt paraphrasings, our findings provide insights into cross-lingual and ethical biases of LLMs and their intersection. We discover significant variance in alignment across languages, challenging the assumption of uniform moral reasoning in AI systems and highlighting the importance of incorporating diverse perspectives in AI ethics. The results underscore the need for further research on the integration of multilingual dimensions in responsible AI research to ensure fair and equitable AI interactions worldwide. [^1] Our code and data are at https://github.com/causalNLP/multiTP.
@@ 1 Introduction
The increasingly impressive performance of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Bubeck et al., 2023) also brings safety concerns. First, do LLMs align with human preferences Bengio et al. (2023); Hendrycks et al. (2021b); Anwar et al. (2024)? Second, whose preferences are they most aligned with (Sorensen et al., 2024c)? Prior work on evaluating alignment often takes one of two forms. The first requires continuous real-time evaluation, such as in ChatArena Chiang et al. (2024), where users select which of two responses they prefer. The second is traditional benchmarking with a static dataset where the answers are determined by the judgment of an experimenter or crowd-sourced from a limited online panel Sap et al. (2019); Forbes et al. (2020); Atari et al. (2023); Jiang et al. (2021b).
@@
The MultiTP dataset includes 98,440 trolley problem scenarios, on which we evaluate 19 different LLMs (open and closed-weight). Our investigation is structured by five research questions, ranging from measuring the overall alignment across all scenarios to whether (mis)alignment correlates with whether a language is high- or low- resource. Our findings reveal that very few LLMs demonstrate overall alignment with human preferences. Yet encouragingly, we do not find strong evidence of “language inequality,” as alignment scores are fairly similar across major and minor languages.
Main contributions. This work (1) presents Multilingual Trolley Problems, a test set enabling thorough inspection of LLM multi-lingual alignment with human preferences, (2) deploys parametric variation in trolley problem queries, to inspect the causal effect of six different preference dimensions, (3) investigates the alignment of 19 LLMs with human preferences through over 100 languages, and (4) finds that most LLMs do not align well with human preferences on trolley problems, but also in the meantime do not demonstrate a bias towards low-resource languages.
@@ 3.1.0.0.1 The Original Human Study
We base our study on the extensive set of human preferences collected in Awad et al. (2018), which serves as the ground truth for alignment. This collection includes approximately 40 million responses from individuals in 233 countries. Following the original study’s methodology, we include data from countries with at least 100 respondents, resulting in a dataset that encompasses 130 countries. Note that while this test set serves as a descriptive measure of human responses to moral dilemmas, it does not prescribe normative guidelines for how systems should act, as the dataset reflects culturally specific biases that may not align with universally accepted ethical principles.
@@ 3.2.0.0.2 Multilingual Variation
We initially developed 460 English vignettes by systematically varying scenarios and character combinations across six moral dimensions, ensuring thorough coverage of all possible character interactions. To expand the scope of our study, we translated this dataset into multiple languages. Using the googletrans Python package, we employed Google Translate to convert the English prompts into 107 supported languages. Although this does not cover every language globally, it encompasses a wide variety, including many low-resource languages with relatively few speakers. Instead of using LLMs, whose translation quality is still uncertain, we use Google Translate, as it is widely recognized as a reliable tool for translating English into other languages, especially for less common languages (Costa-jussà et al., 2022; Jiao et al., 2023; Zhu et al., 2023; Peng et al., 2023). A full list of the languages used in our trolley problem scenarios is provided in Section B.1. To ensure the accuracy of these translations, we manually reviewed a subset of them in several major languages to confirm that the intended meaning of the prompts was preserved. We conducted an Amazon Mechanical Turk task to evaluate translation quality across 44 languages, with detailed information about the evaluation provided in Section D.1.
@@ 3.2.0.0.3 Dataset Statistics
We present comprehensive statistics of the MultiTP dataset in Table 2. The dataset comprises 98,440 trolley problem vignettes, with 460 vignettes for each of the 107 languages. We also provide a snapshot of statistics for four representative languages from different global regions: global west, east, and south. On average, vignettes contain 51 words; English scenarios average 47 words, while Chinese scenarios, due to the linguistic structure, average 78 words per scenario. In contrast, Swahili vignettes are shorter, averaging 38 words. The entire dataset incorporates 6,492 unique words, with a type-token ratio of 0.0013.
@@ 4.0.0.0.2 Preference Assessment
Our test is designed to include six types of systematic variations, allowing us to represent a model’s moral preference with six values: $p_{\text{species}},p_{\text{gender}},p_{\text{fitness}},p_{\text{status}},p_{\text{age}},p_{\text{number}}$. For each dimension $p_{i}$, we report the percentage $p_{i}\in[0,1]$ of the time when a default value prevails, namely sparing humans (over pets), sparing more lives (over fewer lives), sparing women (over men), sparing the young (over the elderly), sparing the fit (over the less fit), and sparing those with higher social status (over lower social status). For example, if a model’s $p_{\text{species}}=1$, it consistently prefers humans over pets, 100% of the time.
@@ 4.0.0.0.3 Misalignment Metric
Calculating misalignment presents a significant challenge due to the different bases of recording human and LLM preferences — by country for humans and by language for LLMs. While in many instances a language corresponds to a country, such as Italian to Italy and Romanian to Romania, there are numerous countries where multiple languages are spoken. In these cases, we compute a weighted average of the misalignment scores for all languages spoken within the country, using the number of speakers per language as weights. We source our language population statistics from Wikipedia and detail our method for mapping languages to countries in Section B.2. We also discuss the potential limitations of this approach in Section 7.
@@ 5 Research Questions and Results
Our study explores five research questions (RQs) that examine various aspects of the moral alignment of LLMs with human preferences. We begin by assessing the global alignment of LLM preferences with human preferences (RQ1). Next, we analyze how LLMs respond to the six principal dimensions studied in the Moral Machine trolley problems (RQ2). We then investigate if LLMs’ responses vary significantly across different languages and identify clusters of language groups where LLMs exhibit similar behavior (RQ3). We also test a “language inequality” hypothesis to determine if LLMs are more likely to be aligned with high-resource language than with low-resource languages (RQ4). Finally, we conduct a robustness study to evaluate the consistency of LLM responses to various paraphrasings of the same trolley problem prompt (RQ5).
@@ 5.1 RQ1: Do LLMs Align with Human Preferences Overall?
(a) Human preferences.
(b) Llama 3.1 70B (aligned).
(c) GPT-4o Mini (misaligned).
@@ 5.2 RQ2: What Are LLMs’ Preferences on Each Moral Dimension in Multilingual Trolley Problems?
To further understand how each moral dimension contributes to the overall misalignment, we find strong correlations between the overall misalignment and several moral dimensions, especially gender, age, and fitness. With a p-value of less than 0.001, we find the correlation with gender is 0.87, 0.69 for age, and 0.68 for fitness. In Figure 2(b), we show the distribution of preference scores across these dimensions for both the three most aligned models (in green) and the three least models (in red). This analysis confirms that misaligned models often exhibit extreme preferences on each moral dimension, such as a higher propensity to protect females, the young, and individuals of higher social status, diverging markedly from the more balanced tendencies observed in human subjects.
Table 3: Language sensitivity scores for all models. Higher scores mean more varied misalignment across languages. We highlight the models with the most and least language sensitivity scores.
Model Sensitivity
@@
Figure 4: Distribution of preferences by each moral dimension across languages, using the most-aligned model Llama 3.1 70B. The dashed line is the overall human preference on each dimension. Cluster A: Georgian, Filipino, Maltese, etc. B: German, Italian, Ukrainian, etc. C: English, Finnish, Chinese, etc. D: Hungarian, Kazakh, Uyghur, etc. See Section E.1 for the entire list of languages in each cluster, as well as clustering results for other models such as GPT-3 and GPT-4.
@@ 5.3 RQ3: Does LLMs Behavior Depend on the Language of Trolley Problems?
The language sensitivity scores of 19 LLMs, as documented in Figure 4, show that most models exhibit clear sensitivity across their language-specific responses, with standard deviation values ranging from 14.7 to 24.7 across languages. Exploring further, we identify distinctive patterns within the language clusters using K-Means clustering. In Llama 3.1 70B, we identify four distinct language clusters. For example, Cluster A (e.g., Georgian and Filipino) values animal lives more than the other three clusters, which unanimously always favor humans over animals. Cluster D (e.g., Hungarian and Kazakh) has a close-to-zero bias against unfit people, with a 50% possibility favoring the fit over the unfit.
@@ 5.4 RQ4: Are LLMs More Misaligned in Low-Resource Languages?
A considerable amount of recent research indicates that LLMs tend to align more closely with high-resource languages, as noted in various studies (Tanwar et al., 2023; Cahyawijaya et al., 2023; Nguyen et al., 2024; Ryan et al., 2024; Zhao et al., 2024). Despite this trend, we seek to determine if the same alignment preference applies to multilingual trolley problems. An important property of MultiTP is that it uses data that is not found online and hence was not seen during the LLMs’ training. This is because we generate our test data on-demand based on parametric variations and translations that we perform specifically for this study. In this context, our research question investigates whether there is a significant positive correlation between the degree of LLM alignment on trolley problems and how widely used a language is. We measure the number of speakers per language by extracting data from Wikipedia’s language demographics statistics Wikipedia (2024). Then, we calculate the Pearson correlation coefficient between model misalignment scores and the number of speakers for each language.
@@
Most models do not show a significant correlation between the per-language misalignment score and the number of speakers per language, with correlations close to zero. See the entire table of the correlation coefficients of all models in Section E.2. This is an interesting observation unique to the MultiTP trolley problems, which challenges the necessity of the “language inequality” hypothesis. For example, in Llama 3.1 70B, the misalignment scores for the top five most spoken languages are: Chinese 0.38, Hindi 0.51, English 0.58, Spanish 0.68, and Arabic 0.54. This distribution of correlations is about the same as in less spoken languages among our dataset: Bosnian 0.54, Luxembourgish 0.37, Icelandic 0.58, Maltese 0.57, Malayalam 0.54, and Catalan 0.59. We also visualize a misalignment world map in Section E.4 showcasing that there is no strong correspondence between the development level of a country and its alignment score.
@@ 5.5 RQ5: Are LLMs Robust To Prompt Paraphrases?
6 Jailbreaking LLMs to Reduce Refusal Rates
Exploring whether jailbroken models maintain similar moral preferences is important, particularly to measure increased bias towards gender or social status. Jailbreaking also helps circumvent LLM refusals. Due to safety tuning in recent models, a key challenge in our study is the high refusal rate; LLMs often decline decisions involving human lives or give generic responses. Earlier models like GPT-3 have a relatively low refusal rate (12.1%), but this rises significantly in newer models. When addressing moral dimensions, models provide definitive answers on species and lives saved but avoid sensitive topics like gender. We employed a basic token-forcing technique to reduce refusals.
Beyond this, we explore advanced jailbreaking methods to further lower refusal rates and reveal underlying preferences. Following Arditi et al. (2025), we apply an uncensoring technique to models with high refusal rates, steering them away from refusals on ethically sensitive instructions. Our experiments on 4 open-sourced LLMs, shown in Figure 2(a) and Appendix F, indicate that uncensored models better align with human responses and reduce refusals across all six dimensions. However, refusal rates do not drop to zero, highlighting areas for future jailbreaking research.
@@ 7 Future Work Directions
Beyond the jailbreaking experiment and the research questions we have asked and answered above, we highlight several areas for improvement and productive future study.
Extension to Different Modalities and More Variations: In our study, we focus exclusively on a text-only setting to investigate the latest text-based LLMs available at the time of our research. Although the original Moral Machine study included a visual demonstration of each scenario, [^2] https://moralmachine.net we recommend that future research also test the alignment of multi-modal foundation models. Additionally, there is an opportunity for future studies to expand the diversity of characters, scenario setups, and dimensions of variation. This expansion would increase the combinatorial variation of moral dilemmas that can be posed as trolley problems. Another point to consider is that the trolley problem is sometimes criticized for being too narrow or unrealistic Steen (2024). Nonetheless, it remains a widely used tool for studying moral decision-making. For this reason, the current study adopts this foundational scenario while encouraging future research to incorporate additional complexities and connect them to real-world situations.
More Language Support: Human languages are vast and diverse. While our study utilizes the 107 languages supported by high-quality translation services from Google Translate, there is a need for further research into low-resource languages. We recommend that future studies expand their scope to include a broader range of these low-resource languages to gain a more comprehensive understanding of LLMs’ moral preferences in diverse linguistic contexts.